A company has $20 million in assets and a total asset turnover ratio of 2. Its costs are equal to 30% of sales. The firm has ROE of 20% and a NPM of 8%. Assume the firm doesn’t have any preferred stock. If the firm has $2 million in cash and Market Value of Equity to Book Value of Equity (M/B) is 3.5, what is its EV/EBITDA?

Now assume that the industry average for EV/EBITDA is 3. If the firm has 2 million shares outstanding, what is the expected stock price per share? Hint: Use the industry average to work towards an expected equity amount. Then divide E(Equity) by shares outstanding.
Answer #1

2 - Net sales 20 2x 20 = Net sales Net sales = 40 million

Cost 30 % of sales = 12 million EBITDA = SALES - TOTAL COST = 40 - 12 = 28 million

sales Net profit margin-Net profit 0.08 - Net profit 40 0.08 x 40 = Net profit Net profit = 3.2 million

net profit Return on equity= 2 share holders equity Share holders equity = 3.270.2 = 16

Market value Market value of equity to book value of equity = book value 35 - Market value 16 Market value = 3.5x16 Market va

EV = Market capitalisation+ Preferred share +Minority interest +Debt -cash and cash equivalents.

Here we have EV = 56 +0+0 +0 -2 =54 million

EBITDA = earnings before interest, taxes, depreciation and amortisation.

EBITDA =28 Million

EV EV / EBITDA= EBITDA 54 EV / EBITDA= 28 EV | EBITDA = 1.92 or 2 after rounding off